# Can Input Attributions Interpret the Inductive Reasoning Process Elicited in In-Context Learning?

### v1

@@ 4.2 Baselines
We also examine a somewhat naive approach of directly asking the models to generate their rationale. Specifically, we have models generate the most informative example in a prompt (Appendix C) in deriving their answer.
@@ 5 Experiments
Figure 2 shows the results in 10-shot settings. [^4] Additional results using fine-tuned models and non-contrastive IA methods are provided in Appendix D.
@@ 5.1 Main results
Input erasure works well
The IE method yielded the best attribution accuracy in identifying the most informative example in most cases. However, it is computationally costly since it requires multiple forward passes with different inputs ($O(n^{3})$ compared to the standard $O(n^{2})$ complexity). That is, if the computational costs are not the problem, simply applying the input erasure is the most reliable way to estimate the importance of each example in ICL. At the same time, despite our controlled task settings and nearly 100% accuracy of LLMs in yielding the correct answer, the IE method does not yield the perfect attribution accuracy in some tasks; this limitation is discussed in § 6. Notably, baselines, such as edit distance of the examples, did not yield a good attribution accuracy overall, which proves that our test suites are challenging.
Improved versions of gradient norm do not show advantage
Among the gradient-based methods, simple gradient norm tends to work the best in most (i.e., LD, AM, VO, SP, ST) tasks. In other words, whereas I $\times$ G and IG are proposed as the improved version of simple gradient norm method, there were no substantial advantages of these methods in our settings. In particular, IG consistently yielded the lowest attribution accuracy across all six tasks among the gradient-based methods, suggesting its limitations in ICL scenarios. The plausible reason behind this inferiority is discussed in § 6.
@@ 5.2 Scaling properties
LLMs have progressively been scaled up toward model parameter size and its context length in the NLP community. We analyze how such a scaling affects the LLMs’ interpretability, particularly in our setting.
@@
We first investigate the relationship between attribution accuracy and model size — is it more difficult to interpret larger LLMs? We observe somewhat intriguing patterns for this question (Figure 5); gradient-based methods tend to work worse in larger LLMs, and in contrast, the self-answer baseline works better in larger LLMs. That is, the (empirically) appropriate approach to interpret the LLMs may differ in their model scale.
Figure 3: Attribution accuracy across six tasks for different Gemma-2 model sizes. Results demonstrate that gradient-based attribution methods show declining performance in larger LLMs, while the self-answer exhibits improved performance with increasing model size.
Figure 4: Attribution accuracy for interpreting Gemma-2-2B models across all six tasks. Gradient-based methods is relatively robust to the number of few-shot examples, while there is a consistent, large drop in attribution accuracy in SA. Note that both x-axis and y-axis are in log scale, and we did not conduct experiments on 100-shot due to computation resource limitations.
@@
Next, given the trend of long-context LLMs, we examine the relationship between attribution accuracy and the number of few-shot examples. Figure 4 shows the attribution accuracy for Gemma-2-2B across all six tasks. This demonstrates that gradient-based methods maintain accuracy or rather improve against the longer context, in contrast to the decreasing chance rate. [^5] For IG in this setting, due to computation resource limitations, we did not conduct experiments. This suggests the robustness of IA methods in long-context scenarios, highlighting their potential for interpreting inputs with extensive contextual information. Notably, the quality of self-answer consistently degraded as the number of in-context examples increased, while the positive scaling effect observed in the previous analysis of model size. That is, the gradient-based methods and self-answer approach exhibit an insightful trade-off between different scaling properties.
Task Acc. (%) IE Worked?
LD (a) $98.0$ No
LD (b) $0.5$
AM (a) $34.5$ Yes
AM (b) $65.5$
VO (a) $100.0$ No
VO (b) $0.0$
SP (a) $98.0$ No
SP (b) $2.0$
ST (a) $68.5$ No
ST (b) $31.5$
@@ 6 Discussion
Why did IE fail to ahieve 100% attribution accuracy?
Our tasks can not be answered without disambiguating examples. Thus, it is somewhat unintuitive to see the non-100% attribution accuracy of the IE method (again, the LLMs understood the task as they achieved 100% accuracy in the tasks) — what happens here? To obtain a hint to clarify IE’s potential limitations, we analyze model behaviors when the disambiguating example is excluded. Interestingly, LLMs adopted a specific generalization (rule) in each task when there was no disambiguating example (Table 2); in other words, they sometimes exhibited strong inductive biases in our tasks. That is, when the correct rule is equal to their preferred rule by their inductive bias, they can answer the task correctly even without disambiguating examples, and the IE method does not compute a proper attribution score. It is now common to see LLMs have particular inductive biases (not a tabula rasaWarstadt et al. (2020); Kharitonov and Chaabouni (2021). Catching such generalization bias with IA methods represents an inherent challenge, highlighting their potential limitations in interpreting ICL.
@@ 7 Related Work
Several lines of research are conducted to interpret neural language models. NLP researchers have adapted IA methods, which was originally applied to vision models Simonyan et al. (2014); Springenberg et al. (2015); Zintgraf et al. (2017), to perform a post-hoc interpretation of input-output associations exploited by language models Karpathy et al. (2015); Li et al. (2016a); Arras et al. (2016); Lei et al. (2016); Alvarez-Melis and Jaakkola (2017), and its improved versions have also been developed Denil et al. (2014); Sundararajan et al. (2017); Murdoch et al. (2018); Sinha et al. (2021); Ding and Koehn (2021); Bastings et al. (2022); Yin and Neubig (2022); Ferrando et al. (2023). In line with these studies, we provide a new perspective to evaluate these IA methods in ICL. Note that, as an orthogonal attempt, some research estimates the saliency scores to directly prompt models to generate such explanations Rajani et al. (2019); Liu et al. (2019); Wu and Mooney (2019); Narang et al. (2020); Marasovic et al. (2022). This method is indeed examined as one baseline in our study.
@@ 8 Conclusions
Regarding model selection, we concentrated on widely-used open-weight LLMs. Since applying IA methods requires gradient computation through backward propagation, computational constraints limited our ability to evaluate all available models, particularly large ones such as Llama-2-70B Touvron et al. (2023) and Llama-3.1-405B Dubey et al. (2024), and limited us to evaluate IG in § 5.2.
@@
We focused exclusively on pre-training models, excluding post-training models from our analysis. This choice was motivated by our interest in basic few-shot learning, which is more commonly used with pre-trained models. Although post-training models might demonstrate higher accuracy on our tasks and the self-answer setting due to their potentially superior capabilities Rivière et al. (2024), our primary focus was on evaluating IA methods rather than model performance.
@@
Acknowledgements
We would like to express our gratitude to the members of the Tohoku NLP Group for their insightful comments. And special thanks to Sho Yokoi for his valuable suggestions on how to improve clarity in several aspects. This work was supported by the JSPS KAKENHI Grant Number JP19H04162, JP22J21492; JST Moonshot R&D Grant Number JPMJMS2011 (fundamental research); JST BOOST, Japan Grant Number JPMJBS2421.
@@ Appendix A Finetuning Details
Fine-tuning data consisted of equal numbers of 10-shot, 50-shot, and 100-shot examples.
For each task, we create the train set for fine-tuning use tokens with no overlaping with the test set (the dataset we used for our main experments).
We fine-tuned models on each task seperately, this allowes them to focus solely on learning the underlying rules of that task, without needing to identify which task is being presented during testing
This results in six fine-tuned models per LLM, one for each task, with the exceprtion of Gemma-2-27B model, which we didn’t fine-tune Gemma-2-27B on the Associative-Recall task as the original model performs good enough on this simple task.
@@ A.1 Finetuing parameters
We use a fixed LoRA configuration with rank $r=32$ and scaling factor $\alpha=64$, applying a dropout rate of $0.05$ across all linear modules. The LoRA adaptation includes bias terms in the training.
For optimization, we perform a learning rate sweep using a cosine scheduler with a $5\%$ warm-up period relative to the total training steps. The optimal learning rate typically falls in the order of $1\times 10^{-5}$ when the loss reaches its minimum.
In our experiments, the Llama-2 models achieve zero loss, which is expected in such synthetic setting. The Gemma-2 models, however, converge to final loss values of approximately $0.2$.
A.2 Zero-shot Question Accuracy after Finetuning
We evaluate performance using exact match accuracy, with results presented in Tables 3456, and 7. In the zero-shot setting, some tasks show accuracies significantly below chance rate (indicated in parentheses), as models occasionally generate responses outside the expected answer space. Notably, all models achieve zero-shot accuracies at or below chance rate across all tasks, suggesting that no superficial patterns were learned during fine-tuning to solve these tasks.
@@ Appendix B Base Model Results
Figure 5 presents the attribution accuracies of the base models. While the overall attribution accuracies for linguistic pattern tasks are relatively low, the performance trends across different tasks and models exhibit similar patterns to those observed in the fine-tuned models (Figure 2).
Appendix C Prompt
We present sample of the exact prompt we used for our task, including the ones we used for testing attribution accurices and modified prompts for self-answer.
@@ Appendix C Prompt
Among the 10 examples labeled <0> to <9>, select the single most helpful example for determining the answer to the <target> question. The correct answer to the target question is “6". To conclude this answer, we need to find one example that provides the necessary information. Therefore, the most helpful example is <
Appendix D Non-contrastive IA Methods Results
We further evaluate attribution accuracies using non-contrastive IA methods (specifically, non-contrastive GN, non-contrastive-I×$\times$×G, and non-contrastive IE), as shown in Figure 6. The results exhibit patterns that are highly consistent with those obtained from contrastive methods (Figure 2).
Appendix E More Results on Interpretability vs. Few-shot Number
Table 8910, and 11 show the Pearson correalation for Gemma-2-9B, Gemma-2-27B, Llama-2-7B and Llama-2-13B.
@@ Appendix D Non-contrastive IA Methods Results
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $-0.39$ $0.05$ $0.44$ $0.05$ $-0.93$ $0.16$
LD $1.00$ $0.35$ $0.89$ $0.07$ $-0.88$ $0.09$
AM $0.88$ $0.27$ $1.00$ $0.02$ $-0.89$ $0.08$
VO $0.99$ $0.13$ $-0.96$ $0.15$ $-0.88$ $0.10$
TA $-0.90$ $0.01$ $-0.73$ $0.11$ $-0.90$ $0.09$
PT $-1.00$ $0.07$ $-0.90$ $0.28$ $-0.93$ $0.09$
Table 8: TThe Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Gemma-2-2B.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $-0.07$ $0.07$ $1.00$ $0.41$ $-0.86$ $0.90$
LD $-0.75$ $0.07$ $0.00$ $0.04$ $-0.85$ $0.33$
AM $0.98$ $0.09$ $-0.93$ $0.02$ $-0.88$ $0.06$
VO $0.76$ $0.14$ $0.33$ $0.05$ $-0.85$ $0.14$
TA $0.86$ $0.36$ $-0.93$ $0.18$ $-0.90$ $0.09$
PT $0.90$ $0.22$ $0.65$ $0.06$ $-0.86$ $0.07$
Table 9: The Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Gemma-2-9B.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $0.28$ $0.06$ $-0.73$ $0.11$ $-0.86$ $0.29$
LD $0.97$ $0.24$ $-0.69$ $0.08$ $-0.89$ $0.09$
AM $0.98$ $0.26$ $-0.07$ $0.03$ $-0.85$ $0.37$
VO $-0.62$ $0.07$ $0.84$ $0.12$ $-0.86$ $0.15$
TA $0.96$ $0.19$ $-0.48$ $0.06$ $-0.92$ $0.10$
PT $0.90$ $0.14$ $0.25$ $0.04$ $-0.89$ $0.09$
Table 10: The Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Llama-2-7B.
Task GN I×$\times$×G SA
Corr. $\Delta$ Corr. $\Delta$ Corr. $\Delta$
AR $-0.74$ $0.06$ $0.32$ $0.07$ $-0.89$ $0.49$
LD $1.00$ $0.11$ $-0.85$ $0.07$ $-0.87$ $0.14$
AM $1.00$ $0.49$ $-0.97$ $0.14$ $-0.87$ $0.15$
VO $-0.81$ $0.04$ $-0.70$ $0.02$ $-0.84$ $0.43$
TA $0.99$ $0.13$ $0.98$ $0.04$ $-0.87$ $0.44$
PT $0.99$ $0.14$ $-0.87$ $0.02$ $-0.86$ $0.18$
Table 11: The Pearson correlation coefficient between attribution accuracy and the number of few-shot examples for Llama-2-13B.
Figure 5: Attribution accuracies for each task for base models.
Figure 6: Attribution accuracy using non-contrastive methods on fine-tuned models.

### v2

@@ 4.3 Baselines
We also examine directly asking the models to generate their rationale. Specifically, we have models generate the most informative example in a prompt (Appendix D) in deriving their answer as a post-hoc explanation. This might be, more or less, relevant to the verbalization of aha moment recently observed in DeepSeek models Guo et al. (2025).
@@ 5 Experiments
Figure 2 shows the results in 10-shot settings, with both top-2 and top-1 metrics. Additional analyses are presented in Appendices CE, and F.
@@ 5.1 Main results
IE works the best
First of all, the input erasure method generally performed the best in both top-1 and top-2 accuracies. This is somewhat obvious because our task is designed to be unsolvable by removing the aha example and thus rather serves as a quick check for our experimental design. Having said that, the input erasure method has some disadvantages in regard to the computational costs of repeated decoding by removing examples one by one as well as the unclarity of by which unit an erasure should be applied, especially in a real, somewhat noisy input. Additionally, the accuracy was not 100% in almost all the cases; we further discuss the potential flaw of this approach in § 6.
Potential of gradient-based approaches
As for baselines, while the self-answer approach worked well in specific settings (associative recall with larger models), most baselines, including attention weights, generally failed to achieve high accuracy. Edit distance was a somewhat strong baseline, but it has obvious limitations of lacking semantic similarities and was frequently outperformed by GN. Compared to such baselines, the gradient-based method worked relatively well, highlighting the potential of this direction.
Improved versions of gradient-based methods do not outperform GN
Among the gradient-based methods, simple gradient norm tends to work the best in most tasks, especially in top-2 accuracy. In other words, whereas I×$\times$×G and IG are proposed as the improved version of simple gradient norm method, there were no substantial advantages of these methods in our settings. In particular, IG consistently yielded the lowest attribution accuracy across all six tasks among the gradient-based methods, suggesting its limitations in ICL scenarios. The plausible reason behind this inferiority is discussed in § 6.
General failure
Nevertheless, some simple tasks, such as Verb-Object, Tense-Article, and Pos-Title, were ever hard to interpret with any approach. This opens a new field for developing a better interpretation method for ICL.
@@ 5.2 Scaling properties
In the age of LLMs, the setting has progressively been scaled up toward model parameter size and context length. We analyze how such a scaling affects the LLMs’ interpretability.
@@
We first investigate the relationship between attribution accuracy and model size — is it more difficult to interpret larger models? We observe somewhat intriguing patterns for this question (Figure 2); gradient-based methods tend to work worse in larger models, and in contrast, the self-answer baseline works better in larger models (especially in Linear-or-Distinct and Associative-Recall). That is, the (empirically) accurate approach to interpreting the LLMs may differ in their model scale, and the success in interpreting smaller models does not always entail the success in interpreting larger models.
Figure 3: Attribution accuracy for interpreting Gemma-2-2B models across all six tasks. Gradient-based methods are relatively robust to the number of few-shot examples, while there is a consistent, large drop in attribution accuracy in SA. Note that both x-axis and y-axis are in log scale.
@@
Next, given the trend of long-context LLMs, we examine the relationship between attribution accuracy and the number of few-shot examples. Figure 3 shows the attribution accuracy for Gemma-2-2B across all six tasks in different number of in-context examples. This demonstrates that gradient-based methods maintain accuracy or rather improve against the longer context, in contrast to the decreasing chance rate. This suggests the robustness of IA methods in long-context scenarios, highlighting their potential for interpreting inputs with extensive contextual information. Notably, the quality of self-answer consistently degraded as the number of in-context examples increased, while the positive scaling effect was observed in the previous analysis of model size. That is, the gradient-based methods and self-answer approach exhibit an insightful trade-off between different scaling properties.
Task Accuracy (%) IE Attr. Acc. (%)
Top-2 Top-1
LD (Rule A) $98.0$ $58.2$ $56.0$
LD (Rule B) $0.5$
AM (Rule A) $34.5$ $93.1$ $92.2$
AM (Rule B) $65.5$
VO (Rule A) $100.0$ $56.1$ $50.3$
VO (Rule B) $0.0$
SP (Rule A) $98.0$ $52.5$ $50.0$
SP (Rule B) $2.0$
ST (Rule A) $68.5$ $59.0$ $51.1$
ST (Rule B) $31.5$
@@ 6 Discussion
Why did IE fail to achieve 100% attribution accuracy?
Our tasks can not be answered without disambiguating aha examples. Thus, it is somewhat unintuitive to see the non-100% attribution accuracy of the IE method (again, the LLMs understood the task as they achieved 100% accuracy in the tasks) — what happens here? To obtain a hint to clarify IE’s potential limitations, we analyze model behaviors when the disambiguating example is excluded. Interestingly, LLMs adopted a specific generalization (rule) in each task when there was no disambiguating example (Table 2); in other words, they sometimes exhibited strong inductive biases in our tasks. That is, when the correct rule is equal to their preferred rule by their inductive bias, they can answer the task correctly even without disambiguating examples, and the IE method does not compute a proper attribution score. It is now common to see LLMs have particular inductive biases (not a tabula rasaWarstadt et al. (2020b); Kharitonov and Chaabouni (2021). Catching such generalization bias with IA methods represents an inherent challenge, highlighting their potential limitations in interpreting ICL.
@@ 7 Related Work
Several lines of research are conducted to interpret neural language models. NLP researchers have adapted IA methods, which were originally applied to vision models Simonyan et al. (2014); Springenberg et al. (2015); Zintgraf et al. (2017), to perform a post-hoc interpretation of input-output associations exploited by language models Karpathy et al. (2015); Li et al. (2016a); Arras et al. (2016); Lei et al. (2016); Alvarez-Melis and Jaakkola (2017), and its improved versions have also been developed Denil et al. (2014); Sundararajan et al. (2017); Murdoch et al. (2018); Sinha et al. (2021); Ding and Koehn (2021); Bastings et al. (2022); Yin and Neubig (2022); Ferrando et al. (2023). In line with these studies, we provide a new perspective to evaluate these IA methods in ICL. Note that, as an orthogonal attempt, some research estimates the saliency scores to directly prompt models to generate such explanations Rajani et al. (2019); Liu et al. (2019); Wu and Mooney (2019); Narang et al. (2020); Marasovic et al. (2022). This method is indeed examined as one baseline in our study.
@@ 8 Conclusions
Regarding model selection, we concentrated on widely-used open-weight LLMs. Since applying IA methods requires gradient computation through backward propagation, computational constraints limited our ability to evaluate all available models, particularly large ones such as Llama-2-70B Touvron et al. (2023).
@@
We focused exclusively on pre-training models, excluding post-training models from our analysis. This choice was motivated by our interest in basic few-shot learning, which is more commonly used with pre-trained models. Although post-training models might demonstrate higher accuracy on our tasks and the self-answer setting due to their potentially superior capabilities Riviere et al. (2024), our primary focus was on evaluating IA methods rather than model performance.
@@ Appendix A Finetuning details
The fine-tuning dataset consisted of 400 tasks for each of the 10-shot, 50-shot, and 100-shot settings (1,200 tasks total). For each task, we created a training set for fine-tuning using tokens that did not overlap with our test set (the dataset used in our main experiments). We fine-tuned models separately on each task, resulting in six fine-tuned models per LLM. The exception was the Gemma-2-27B model, which we did not fine-tune on the Associative-Recall task since the original model already performed well enough on this simple task.
@@ A.1 Finetuing parameters
We use a consistent LoRA configuration with rank $r=32$ and scaling factor $\alpha=64$, applying a dropout rate of $0.05$ across all linear modules. The LoRA adaptation includes bias terms in the training. For optimization, we perform a learning rate sweep using a cosine scheduler with a $5\%$ warm-up period relative to the total training steps. The optimal learning rate typically falls in the order of $1\times 10^{-5}$ when the loss reaches its minimum. In our experiments, the Llama-2 models achieved nearly zero loss, which is expected in such a synthetic setting. The Gemma-2 models, however, converge to final loss values of approximately $0.2$.
A.2 Zero-shot task accuracy after finetuning
We evaluate task accuracy using exact match, with results presented in Tables 4567, and 8. In the zero-shot setting, some tasks show accuracies significantly below chance rate (indicated in parentheses), as models occasionally generate unexpected responses. Notably, all models achieve zero-shot accuracies at or below the chance rate across all tasks, suggesting that models cannot solve our tasks relying only on the aha example.
@@ Appendix B Base model results
Figure 4 presents the IA scores for the base models. While the overall IA scores for VO, TA and PT tasks are relatively low, the performance trends across different tasks and models exhibit similar patterns to those observed in the fine-tuned models (Figure 2). Therefore, our results can be generalized regardless of fine-tuning.
Appendix C Aggregating attritbuion with max
Figure 5 presents the IA scores using maximum aggregation to convert token-level attribution to example-level attribution: $S(e_{i})=\max_{x_{j}\in(X_{i},y_{i})=e_{i}}S(x_{j})$. The overall trend for all the IA methods is consistent with the sum aggregation (Figure 2); thus, our results can be generalized regardless of this design.
Appendix D Prompts
We present sample of the exact prompt we used for our task, including the ones we used for testing attribution accurices and modified prompts for self-answer. Note that in all the experiments, we only used the model outputs with a correct answer. That is why we appended the correct answer in advance to the self-answer prompt to obtain the post-hoc explanation.
@@ Appendix D Prompts
Among the 10 examples labeled <0> to <9>, select the single most helpful example for determining the answer to the <target> question. The correct answer to the target question is “6”. To conclude this answer, we need to find one example that provides the necessary information. Therefore, the most helpful example is <
Appendix E Chain-of-though format
To contextualize our experimental settings with more practical scenarios, we further evaluate attribution accuracies on top of chain-of-thought (CoT) prompting Wei et al. (2022). We use Gemma-2-27B-IT (a post-training version of Gemma-2-27B) instead of the base model to perform a better CoT-style generation and employ the AM task, where the model achieved high accuracy with CoT, as a case study. We only target the last time step to generate the exact answer. We compute the by-example attribution scores the same as our main experiments, but now the attribution scores can be spread over the reasoning chain part as well as in-context examples. Our target is which example is informative to answer the question; thus, the attribution to the chain part is tentatively disregarded. As statistics, we just report how many proportions of attribution scores ([0-100%]) reached the reasoning chain part (denoted as “Chain prop.”).
The results are presented in Table 3. All tested IA methods performed worse in this CoT setting than in the CoT-free settings in the main experiments. Nevertheless, the superiority of GN to other approaches still holds. Note that the Chain prop. substantially differs across IA methods; for example, IE assigns over 80% of the attribution score to the chain. These divergent results also suggest that the conventional IA methods can not easily be applied to modern ICL and CoT settings.
The exact prompt and the reasoning chain generated by the model are provided below [^3] The prompt template is applied since this is a post-training model:
CoT Prompt
{mdframed}
[ backgroundcolor=black!5, topline=false, bottomline=false, rightline=false, leftline=false, innertopmargin=0.5em, innerleftmargin=0.5em, innerbottommargin=0.5em, innerrightmargin=0.5em, ]<start_of_turn> user
Input: saw, 2, Output: saw, 4
Input: start, 2, Output: start, 4
Input: the, 2, Output: the, 4
Input: too, 2, Output: too, 4
Input: round, 2, Output: round, 4
Input: which, 1, Output: which, 3
Input: work, 2, Output: work, 4
Input: get, 2, Output: get, 4
Input: that, 2, Output: that, 4
Input: white, 2, Output: white, 4
Input: I, 3, Output: <ANSWER>
Solve this problem step by step, generate the content of <ANSWER> after “So the answer is”: <end_of_turn>
<start_of_turn> model
Method IA score Aggregation Chain prop. (%)
Top-1 Top-2
IE $11.3$ $17.4$ Sum $82.1$
GN $12.4$ $37.6$ $35.4$
I $\times$ G $14.9$ $29.8$ $23.0$
IG $8.9$ $22.7$ $42.9$
IE $11.3$ $17.4$ Max $82.1$
GN $14.5$ $33.3$ $19.2$
I $\times$ G $9.9$ $31.2$ $23.1$
IG $9.9$ $22.3$ $7.5$
Table 3: IA score for the CoT-prompted AM task. The percentage of attribution scores allocated to the reasoning chain is denoted as Chain prop.
Appendix F Distribution of example with the highest attribution score
Figures 6, 7, 8, 9 and 10 present the distribution of example positions with the highest attribution scores. All IA methods, except for I $\times$ G, possess positional bias for certain tasks, specifically favoring examples either at the beginning or end, aligning with the position bias known to LLMs Liu et al. (2024).
Appendix G Informaiton for responsibility checklist
We utilized software libraries, including the Huggingface toolkit Wolf et al. (2020) (Apache License Version 2.0), Captum package Kokhlikyan et al. (2020) (BSD 3-Clause License) for IG computation, and PyTorch Ansel et al. (2024) (BSD 3-Clause License). These tools were used according to their licenses and intended usage. We used writing assistance tools (including Grammarly) for language error correction only. The computational budget of this work is approximately 600 GPU hours (with A100/H100/H200 machines).
@@ Appendix G Informaiton for responsibility checklist
Figure 4: Attribution accuracies for each task for base models. Similar patterns to those observed in the fine-tuned models (Figure 2) can be obsered.
Figure 5: Attribution accuracies for each task use max aggregation. The overall trend for all IA methods is consistent with sum aggregation (Figure 2)
Figure 6: Distribution of the positional of the example with the highest attribution scores across IA methods (Llama-2-7B model).
Figure 7: Distribution of the position of the example with the highest attribution scores across IA methods (Llama-2-13B model).
Figure 8: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-2B model).
Figure 9: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-9B model).
Figure 10: Distribution of the position of the example with the highest attribution scores across IA methods (Gemma-2-27B model).